OpenStudy (anonymous):

1. put this into slope-intercept form :y-4= -2/3(x+2) 2. and this: y-6= -2/3(x+5)

OpenStudy (anonymous):

i dont need the answer, i need the work

OpenStudy (anonymous):

plus that website is not good because it thinks i need to solve it, and i dont, i just need it in slope-intercept form.

OpenStudy (oaktree):

Okay, let me help you.

OpenStudy (anonymous):

thnxx

OpenStudy (anonymous):

the first equation was y-4= -2/3(x+2) but i already did the first step of course

OpenStudy (oaktree):

Multiplying by 3?

OpenStudy (anonymous):

what?

OpenStudy (oaktree):

Okay, good. You didn't do that.

OpenStudy (oaktree):

Basically all you want to do is add 4 to both sides and then simplify.

OpenStudy (anonymous):



OpenStudy (anonymous):

so i add 4 to 4, and also to -2/3??

OpenStudy (oaktree):

\[y = \frac{ -2 }{ 3 }x + \frac{ 8 }{ 3 }\]should be your final answer.

OpenStudy (oaktree):

Are you familiar with inverse operations? Like, doing the opposite to both sides?

OpenStudy (anonymous):

wait what did u do

OpenStudy (oaktree):

Added.

OpenStudy (anonymous):

can u show me the work?

OpenStudy (anonymous):

ok fine leave... :(

OpenStudy (anonymous):

i need help with the second one